Winner No Yes No ., Yes Person Alcohol Consumption Yes Suppose you are interested in running the a regression for the effect of winning the lottery on alcohol consumption. Your only regressor in the regression is a dummy variable for winning the lottery. What is the coefficient on the dummy variable for winning the lottery? Continuing the previous question, what is the coefficient on the intercept in that regression?*
Answer #1

Person Winner (categorical) Winner(numeric) (x) x^2 x*y Alcohol Consumption (y)

1 No 0 0 0 4

2 Yes 1 1 6 6

3 No 0 0 0 2

4 Yes 1 1 8 8

5 No 0 0 0 3

6 Yes 1 1 7 7

Sum 3 3 21 30

the intercept is (Sum(y)*sum(x^2)-Sum(x)*Sum(xy))/(n*Sum(x^2)-Sum(x)^2) = (30*3-3*21)/(6*3-3^2) = 3.00000

Intercept 3.0000000 Found using Excel intercept formula

Slope 4.0000000 Found using Excel slope formula